OpenStudy (anonymous):

Find the derivative of tan(2x-1)^2

OpenStudy (anonymous):

ok

OpenStudy (raden):

derivative of (2x-1)^2 = 2(2x-1)*2 derivative of tan(2x-1)^2 = sec^2 (2x-1)^2 use "chain rule"

OpenStudy (anonymous):

not product?

OpenStudy (raden):

just multiply of them, gives y' = 2(2x-1)*2* sec^2 (2x-1)^2 y' = 4(2x-1)* sec^2 (2x-1)^2 y' = (8x-4)*sec^2 (2x-1)^2

OpenStudy (anonymous):

I thought it was sec2(8-4)

OpenStudy (anonymous):

sec 2 from tan and then (2x-1)^2 for second part which is chain

OpenStudy (raden):

multiplying process all derivative parts, its called chain rule